OpenStudy (anonymous):

what is the average kinetic energy of 1 mole of gas at 102 degrees celsius?

OpenStudy (anonymous):

\[KE _{ave} = (3/2) R T\] R= 8.314 J/mol K T= 273 + 102 = 375K \[KE _{ave}\] = (3/2) x 8.314 x 375 = 4680 J/mol